# Safe Output Regulation of Coupled Hyperbolic PDE-ODE Systems

### v1

@@ 4.2 Output-feedback safe control
where $\acute{k}_{i}$ are defined in (74), and where $\mathcal{D}_{v}$ denotes the bounds of the disturbances given in Assumption 4. Besides, there also exist conservatism in the choice of $\sigma(t)$, which arises from using known bounds to estimate the lower bound of $h(e(\bar{t}_{0}),\bar{t}_{0})$, denoted as $\underline{h}(e(\bar{t}_{0}),\bar{t}_{0})=\min_{\ell_{1},\ell_{2}\in\mathcal{D},\ell_{3}\in\mathcal{D}_{v}}\mathcal{P}(Y(t_{0}),\ell_{1}[t_{0}],\ell_{2}[t_{0}],\ell_{3}(t_{0}))$, where the function $\mathcal{P}$ is defined in (51). The lower bound $\underline{h}(e(\bar{t}_{0}),\bar{t}_{0})$ is used in place of $h(e(\bar{t}_{0}),\bar{t}_{0})$ in the coefficient of the function $\sigma(t)$ (49) (the sign of $h(e(\bar{t}_{0}),\bar{t}_{0})$ used in judging the case in (49) is known according to Assumption 5). Design the output-feedback safe controller as
$U_{f}=\hat{U}+{\rm sign}(\vartheta(t_{0}))\rho_{e}(t)$ (125)
where $\hat{U}$ is obtained by replacing the unmeasurable distributed states $z(x,t)$, $w(x,t)$, and the external disturbances $v(t)$ in the nominal control law (72) with their estimates, and where the sign is defined by ${\rm sign}(u)=\begin{cases}1&u\geq 0\\ -1&u<0\end{cases}$ and $\rho_{e}(t)$ is given by
$\rho_{e}(t)=\begin{cases}2\max\{\|\Psi(1,y)\|,\|\Phi(1,y)\|,|\acute{G}_{5}+\lambda(1)|,\xi_{e}\}\rho_{e1},\\ 0\leq t<\frac{1}{q_{1}}+\frac{1}{q_{2}};\\ 2\max\{\|\Psi(1,y)\|,\|\Phi(1,y)\|,|\acute{G}_{5}+\lambda(1)|,\xi_{e}\}\rho_{e2},\\ t\geq\frac{1}{q_{1}}+\frac{1}{q_{2}}.\end{cases}$ (126)
The constant $\xi_{e}$ is defined in Assumption 6. The functions $\rho_{e1}$ and $\rho_{e2}$, which are the upper bounds of observer errors, are given by (122), (123) in Lemma 3. As will be seen clearly later, the function $\rho_{e}(t)$ is indeed an estimated upper bound of the error between the state-feedback and output-feedback controllers. The role of ${\rm sign}(\vartheta(t_{0}))\rho_{e}(t)$ in (125) is to tolerate the estimation errors that arise from the unmeasured distributed states and the external disturbances by shrinking the original safe set to a subset that converges to the original safe set as time approaches infinity.
Remark 1.
Even though there is a switch in (126), this discontinuity can be avoided in the practical implementation by picking a function $\rho_{c}=M_{c}e^{-\sigma_{c}t}$ satisfying $\rho_{c}\geq\rho_{e}$, at the cost of increasing the conservative, i.e., shrinking the safe region.
@@
For initial data $w[t_{0}]\in L^{2}(0,1)$, $z[t_{0}]\in L^{2}(0,1)$, $Y(t_{0})\in\mathbb{R}^{n}$, $v(t_{0})\in\mathbb{R}^{n_{v}}$, $\hat{w}[t_{0}]\in L^{2}(0,1)$, $\hat{z}[t_{0}]\in L^{2}(0,1)$, $\hat{v}(t_{0})\in\mathbb{R}^{n_{v}}$ for design parameters $\kappa_{i}$, $i=1,\ldots,n-1$ satisfying (124), the closed-loop system including the plant (1)–(5), the observer (89)–(94), and the controller $U_{f}$ (125) ensures:
1) There exists a unique solution $w[t]\in L^{2}(0,1)$, $z[t]\in L^{2}(0,1)$, $\hat{w}[t]\in L^{2}(0,1)$, $\hat{z}[t]\in L^{2}(0,1)$, $Y(t)\in\mathbb{R}^{n}$, $v(t)\in\mathbb{R}^{n_{v}}$, $\hat{v}(t)\in\mathbb{R}^{n_{v}}$, to the output-feedback closed-loop system;
2) Tracking error $e(t)=y_{1}(t)-r(t)$ is convergent to zero and all states are bounded;
3) Safety is ensured in the sense that:
@@
$-(\acute{G}_{5}+\bar{\lambda}(1))\tilde{v}(t)-\tilde{p}(1,t).$ (127)
Applying the controller (125), we have
$\beta(1,t)={\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))\rho_{e}(t)-\eta_{e}(t).$ (128)
The target system becomes (67)–(70) with (128). Considering the given initial condition and the transformations (62), (63) together with (3.3), (53), as well as the exogenous signal model (12), we have $\beta[t_{0}]\in L^{2}(0,1)$, $\alpha[t_{0}]\in L^{2}(0,1)$ and $v(t)\in\mathbb{R}^{n_{v}}$. We know from Lemma 3, (73), (120) in Assumption 6, and (126), (127) that ${\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))\rho_{e}(t)-\eta_{e}(t)$ is bounded, and then we have $\beta[t]\in L^{2}(0,1)$, $\alpha[t]\in L^{2}(0,1)$, $H(t)\in\mathbb{R}^{n}$. Applying the transformation (88), through the same process in the proof of property 1 in Theorem 1, we have $w[t]\in L^{2}(0,1)$, $z[t]\in L^{2}(0,1)$, $Y(t)\in\mathbb{R}^{n}$. Recalling the results about the observer errors in Lemma 3, and the definition (95), Property 1 is obtained.
2) Recalling (127), (73), and (120) in Assumption 6, applying Lemma 3, we obtain
@@
$\leq$ $\rho_{e}(t)$ (129)
where $\tilde{\Omega}_{e}(t)$ is defined in (121). We know from (126) and Lemma 3 that $\rho_{e}(t)$ is exponentially convergent to zero, and thus ${\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))\rho_{e}(t)-\eta_{e}(t)$ in (128) is exponentially convergent to zero as well recalling (129). Through the same process in the proof of the property 2 in Theorem 1, Property 2 is obtained.
3) The sign of $\vartheta$ is the same as its initial value at $t_{0}$ because of the fact that $\vartheta$ is continuous and not zero in Assumption 6 (recalling (52)). It is obtained from (128), (129) and (70), (67) that $\vartheta\beta(0,t)\geq 0$ for $t\in[\bar{t}_{0},\infty)$. The choice of the design parameters $k_{i}$ in (124) is a subset of the ones (74), which thus ensures $h_{i}>0$ at $t=\bar{t}_{0}$ according to Lemma 1. Considering the system matrix $A_{\rm h}$ defined in (61) being a high-relative-degree CBFs $h_{i}$, we have $h_{i}>0$ for $t\geq\bar{t}_{0}$. Then, through the same proof as the ones of cases a)–c) of the property 3 in Theorem 1, Property 3 is obtained.
@@ 5.1 Model
Applying Riemann transformations
${z}(x,t)$ $={u_{t}}(x,t)-\sqrt{\frac{T_{0}}{\rho}}{u_{x}}(x,t),$ (134)
${w}(x,t)$ $={u_{t}}(x,t)+\sqrt{\frac{T_{0}}{\rho}}{u_{x}}(x,t)$ (135)
and defining a new variable $Y(t)=[y_{1}(t),\dot{y}_{1}(t)]^{T}$, (130)–(133) can be rewritten as
@@
where $q_{1}=q_{2}=\sqrt{\frac{T_{0}}{\rho}},c=\frac{{{-d_{c}}}}{2\rho},p=q=1$, $A=\left[0,1;0,\frac{-d_{0}}{M_{0}}-\frac{\sqrt{T_{0}\rho}}{M_{0}}\right],B=[0,\frac{\sqrt{T_{0}\rho}}{M_{0}}]^{T},C=[0,2].$ After the transformation (134), (135), the airflow disturbances $f_{i}$ in the wave PDEs is given by the external disturbances $d(t)=\bar{P}_{d}v_{d}(t)$ (14) with $\bar{P}_{d}=I_{4\times 4}$ and
$v_{d}=[\sin(0.25t),\cos(0.25t),\sin(0.5t),\cos(0.5t)]$ (141)
in the $2\times 2$ hyperbolic PDEs (136)–(140), where $G_{1}=[g_{1},g_{2}]^{T}$, $g_{1}=[1,0,0,0]$, $g_{2}=[1,1,1,1]$, $G_{4}=[0,1,0,1]$, $G_{5}=[1,0,1,0]$, $G_{2}(x)=[x,0,0,0]$, $G_{2}(x)=[0,x,0,0]$. The desired reference is set as $r(t)=\sin(0.25\pi t)+\cos(0.25\pi t)$, i.e., $\bar{P}_{r}=[1,1]$, and
$v_{r}=[\sin(0.25\pi t),\cos(0.25\pi t)]$ (142)
in (15). Therefore, the system matrix of exogenous signal model is $S_{d}=[0,0.25,0,0;-0.25,0,0,0;0,0,0,0.5;0,0,$ 0.5,0]-0.5,0] and $S_{r}=[0,0.25\pi;-0.25\pi,0]$.
@@
Compared to the plant described by (1)–(5), the simulation model represented by (136)–(140) includes additional terms, specifically $c_{1}z(x,t)$ and $c_{2}w(x,t)$ in (138) and (139), respectively. However, these additions do not alter the state-feedback control design presented in this paper. To be exact, applying state-feedback control, the target system defined by (67) and (71), which corresponds to the closed-loop system, remains unchanged except for the updates that (69) and (70) are modified to $\beta_{t}=q_{2}\beta_{x}+c\beta$ and $\alpha_{t}=-q_{1}\alpha_{x}+c\alpha$. These modifications clearly do not impact the safety and stability results obtained. The only change in the observer-based output-feedback controller is adding the terms $c\hat{z}(x,t)$ and $c_{2}\hat{w}(x,t)$ in (91) and (92) of the observer, while all other aspects remain unchanged.
The initial values are defined as $w(x,0)=\cos(2\pi x)$, $z(x,0)=\sin(3\pi x)$, $y_{2}(0)=0$, $\hat{w}(x,0)=\cos(2\pi x)+0.5$, $\hat{z}(x,0)=\sin(3\pi x)+0.5$, $\hat{v}_{d}(0)=[0.5,1.5,0.5,1.5]^{T}$, $\hat{v}_{r}(0)=[0.5,1.5]^{T}$ where the initial estimation errors are $0.5$. Especially, the values $y_{1}(t_{0})$ will be given in the next subsections. The simulation is conducted using the finite difference method with the space step as $0.05$ and the time space $0.001$.
@@ 5.2 Controller and simulation results
We test our safe controller in the following two cases. In Case 1, the safe region is on one side of the desired reference of the UAV, i.e., the reference is indeed the barrier; In Case 2, the barriers exist on both sides of the UAV’s desired reference.
5.2.1 Case 1. One side of the desired reference is obstructed.
The barrier function is set as $h(e(t),t)=e(t)=y_{1}(t)-r(t)$, where $r(t)$ is the desired reference of UAV, and where the region above $r(t)$ is safe. Then we know $\vartheta=\frac{\partial h(e(t),t)}{\partial e}=1$. According to (47)–(49), we have $h_{1}=z_{1}(t)+\sigma(t)$, $h_{2}=z_{2}(t)+\dot{\sigma}(t)+k_{1}z_{1}(t)+k_{1}\sigma(t)$, where $\sigma(t)$ is given by (49) with choosing $\epsilon=2$ and $t_{a}=2$. According to (53), $f(\underline{z}_{n}(t),t)$ is obtained as, $f(\underline{z}_{n}(t),t)=\frac{1}{b}((k_{1}+k_{2})z_{2}+k_{1}k_{2}z_{1}+\ddot{\sigma}(t)+(k_{1}+k_{2})\dot{\sigma}(t)+k_{1}k_{2}\sigma(t))$. Then $p(1,t)$ can be obtained by recalling (73) that uses (B.4). Then the state-feedback controller is obtained recalling (72), where the design parameters are chosen as $k_{1}=1.5$, $k_{2}=4$, and the output-feedback controller $U_{f}$ is obtained applying (125) with choosing the control design parameters $k_{1}=3$, $k_{2}=4$, $M_{c}=15$, $\sigma_{c}=2$ recalling Remark 1, and the observer design parameters $L_{d}=[10,8,10,8]$, $L_{r}=[2,1]$. For safe initialization, we set $y_{1}(t_{0})=8$. For the unsafe initialization case, we set $y_{1}(t_{0})=-1$ that is beyond the safe region, and the design parameters in the output-feedback control law are the same as the output-feedback control in the safe initialization except for $k_{1},k_{2}$ that are chosen as $k_{1}=2.5$, $k_{2}=4.5$ now, and additionally choosing $t_{a}=2$, $\epsilon=2$ in (49) to enforce the out return to the safe region. The simulation results are presented below.
@@
(a) Estimates of $\hat{v}_{d}$
(b) Estimates of $\hat{v}_{r}$
Figure 6: Estimates of exogenous signals $v_{d}$, $v_{r}$ in (141), (142) (red lines are estimates and dashed black lines are true values).
Figure 7: Control forces at the top of the cable
We know from the open-loop responses of $y_{1}$ in Fig. 2 that the plant is open-loop unstable, while both state-feedback control and output-feedback control achieve effective safe regulation— tracking the reference and ensuring the position of the payload stays within the safe region. Moreover, when $y_{1}$ starts outside the safe region, the designed control input effectively guides it back to safety within 1.4 seconds, which is less than the designated time $\bar{t}_{0}+t_{a}=2.058$ in the control design. Another ODE state $y_{2}$, i.e., the payload velocity, is shown in Fig. 3, and the PDE states that reflect the oscillations along the cable are given in Fig. 4 (where we only show the case that the output-feedback control from safe initial states for avoiding repetition), demonstrating the boundedness of these signals. We know from Figs 5, 6 that the observer can successfully track the unmeasured PDE states and the exogenous signals $v_{d}$ and $v_{r}$. The control inputs in the above three situations are given in Fig. 7.
@@
Because $h$ is not differentiable at $e=0$, the condition in Assumption 6 is not strictly satisfied by (143). However, the controller remains valid by explicitly defining $\frac{\partial h(e,t)}{\partial e}=-1$ when $e=0$, i.e., $\vartheta(e(t),t)=-1$ for $e(t)>0$ and $\vartheta(e(t),t)=1$ for $e(t)\leq 0$. The signal $\vartheta$ is now discontinuous at $e=0$, violating the continuity requirement in Assumption 6. This implies that ${\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))$ in the output-feedback controller is not guaranteed to be invariant as in the proof of Theorem 2. Nevertheless, with appropriate observer gains $L_{d}$, $L_{r}$ chosen via pole placement, the observer errors can rapidly converge to zero, as shown in Fig., which means that $\rho_{e}(t)$ decays quickly to zero and the influence of the term ${\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))\rho_{e}$ in the output-feedback controller becomes negligible after a short initial period. Therefore, the proposed control design remains valid without modification, as shown in the upcoming simulation results.
According to (47)–(49), we have $h_{1}=-|z_{1}(t)|+\delta(t)+\sigma(t)$, $h_{2}=\vartheta z_{2}+\dot{\delta}(t)+\dot{\sigma}(t)-k_{1}|z_{1}(t)|+k_{1}\delta(t)+k_{1}\sigma(t)$, with $\sigma(t)$ is given by (49) with choosing $\epsilon=2$ and $t_{a}=2$. Recalling (53), we obtain $f$ as $f(\underline{z}_{n}(t),t)=\frac{1}{b}\big((k_{1}+k_{2})\vartheta z_{2}-k_{1}k_{2}|z_{1}(t)|+\ddot{\delta}(t)+(k_{1}+k_{2})\dot{\delta}(t)+k_{1}k_{2}\delta(t)+\ddot{\sigma}(t)+(k_{1}+k_{2})\dot{\sigma}(t)+k_{1}k_{2}\sigma(t)\big)$. Then $p(1,t)$ can be obtained by recalling (73) that uses (B.4). The output-feedback controller $U_{f}$ is obtained applying (125), and recalling (72), with choosing the control design parameters $k_{1}=30$, $k_{2}=16$, $M_{c}=20$, $\sigma_{c}=1$ recalling Remark 1 and the observer design parameters $L_{d}=[10,8,10,8]^{T}$, $L_{r}=[2,1]^{T}$. For safe initialization, we set $y_{1}(t_{0})=-5$. For unsafe initialization, we set $y_{1}(t_{0})=20$, which falls outside the safe region, where the design parameters remain the same as those used for the safe initialization. Additionally, we choose $t_{a}=2$ and $\epsilon=2$ in (49) to ensure that the system’s output returns to the safe region. The simulation results are presented below.
@@
It is observed from Fig 8 that when $y_{1}$, i.e., the position of the payload, starts within the safe region, it successfully tracks the reference meanwhile avoiding collisions with the barriers all the time; when the $y_{1}$ starts outside the safe region, it will return to the safe region at 0.35 s, which is less than $\bar{t}_{0}+t_{a}=2.058$, and also eventually follow the reference. The curves of $y_{2}$ and the PDE states are shown in Figs. 9, 10, which shows that these signals, i.e., the payload velocity and cable oscillations, are bounded, even though anti-damping is added to challenge the task. The results about the observer errors and the estimation errors of the external signals are similar to Figs. 5 and 6, which are not present here for avoiding repetition.

### v2

@@ 4.2 Output-feedback safe control
where $\acute{k}_{i}$ are defined in (73), and where $\mathcal{D}_{v}$ denotes the bounds of the disturbances given in Assumption 5. Besides, there also exists conservatism in the choice of $\sigma(t)$, which arises from using known bounds to estimate the lower bound of $h(e(\bar{t}_{0}),\bar{t}_{0})$, denoted as $\underline{h}(e(\bar{t}_{0}),\bar{t}_{0})=\min_{\ell_{1},\ell_{2}\in\mathcal{D},\ell_{3}\in\mathcal{D}_{v}}\mathcal{P}(Y(t_{0}),\ell_{1}[t_{0}],\ell_{2}[t_{0}],\ell_{3}(t_{0}))$, where the function $\mathcal{P}$ is defined in (51). The lower bound $\underline{h}(e(\bar{t}_{0}),\bar{t}_{0})$ is used in place of $h(e(\bar{t}_{0}),\bar{t}_{0})$ in the coefficient of the function $\sigma(t)$ (49) (the sign of $h(e(\bar{t}_{0}),\bar{t}_{0})$ used in judging the case in (49) is known according to Assumption 6). Design the output-feedback safe controller as
$U_{f}=\hat{U}+{\rm sign}(\vartheta(t_{0}))M_{c}e^{-\sigma_{r}t}$ (137)
where
$\hat{U}=-qz(1,t)+\int_{0}^{1}{\Psi}(1,y)\hat{z}(y,t)dy+\int_{0}^{1}{\Phi}(1,y)\hat{w}(y,t)dy$
$+\lambda(1)\hat{Y}(t)-(\acute{G}_{5}-\bar{\lambda}(1)){\hat{v}}(t)+\hat{\varsigma}(1,t)$ (138)
is obtained by replacing the unmeasurable states $z(x,t)$, $w(x,t)$, $Y(t)$ and and the external disturbances $v(t)$ in the nominal control law (71) with their estimates, and where the sign is defined by ${\rm sign}(u)=\begin{cases}1&u\geq 0\\ -1&u<0\end{cases}$ and $M_{c}$ is given by
$M_{c}=2\max\{\|\Psi(1,y)\|,\|\Phi(1,y)\|,|\acute{G}_{5}+\lambda(1)|,\xi_{e}\}\Upsilon_{2}.$ (139)
The constant $\xi_{e}$ is defined in Assumption 7. The function $\rho_{1e}$ is the upper bound of observer errors, which are given by (132) in Lemma 3. As will be seen clearly later, the function $\rho_{e}(t)$ is indeed an estimated upper bound of the error between the state-feedback and output-feedback controllers. The role of ${\rm sign}(\vartheta(t_{0}))\rho_{e}(t)$ in (137) is to tolerate the estimation errors that arise from the unmeasured states and the external disturbances by shrinking the original safe set to a subset that converges to the original safe set as time approaches infinity.
@@
Assume that the closed-loop system consisting of the plant (1)–(5), the observer (87)–(93), and the controller (137) admits a classical solution. For initial data $w[t_{0}]\in L^{\infty}(0,1)$, $z[t_{0}]\in L^{\infty}(0,1)$, $Y(t_{0})\in\mathbb{R}^{n}$, $v(t_{0})\in\mathbb{R}^{n_{v}}$, $\hat{w}[t_{0}]\in L^{\infty}(0,1)$, $\hat{z}[t_{0}]\in L^{\infty}(0,1)$, $\hat{v}(t_{0})\in\mathbb{R}^{n_{v}}$, and for design parameters $\kappa_{i}$, $i=1,\ldots,n-1$ satisfying (136), the following properties hold:
1) Tracking error $e(t)=y_{1}(t)-r(t)$ is convergent to zero and all states are bounded;
2) Safety is ensured in the sense that:
@@
$-(\acute{G}_{5}+\bar{\lambda}(1))\tilde{v}(t)-\tilde{\varsigma}(1,t),$ (140)
where $\tilde{\varsigma}(1,t)=\varsigma(1,t)-\hat{\varsigma}(1,t)$. Applying the controller (137), we have
$\beta(1,t)={\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))M_{c}e^{-\sigma_{r}t}-\eta_{e}(t).$ (141)
The target system becomes (68)–(69) with (141).
Recalling (139), (140), (72), and (130) in Assumption 7, applying Cauthy-Schwarz inequality and Lemma 3, we obtain
@@
$\leq$ $M_{c}e^{-\sigma_{r}t}$ (142)
where $\tilde{\Omega}_{e}(t)$ is defined in (131). Therefore, $\beta(1,t)$ in (141) is exponentially convergent to zero. Through the same process in the proof of the property 1 in Theorem 1, Property 1 is obtained.
2) The sign of $\vartheta$ is the same as its initial value at $t_{0}$ because of the fact that $\vartheta$ is continuous and not zero in Assumption 7 (recalling (52)). It is obtained from (141), (142) and (69), (68) that $\vartheta\beta(0,t)\geq 0$ for $t\in[\bar{t}_{0},\infty)$. The choice of the design parameters $k_{i}$ in (136) is a subset of the ones (73), which thus ensures $h_{i}>0$ at $t=\bar{t}_{0}$ according to Lemma 1. Considering the system matrix $A_{\rm h}$ defined in (62) being a high-relative-degree CBFs $h_{i}$, we have $h_{i}>0$ for $t\geq\bar{t}_{0}$. Then, through the same proof as the ones of cases a)–c) of the property 3 in Theorem 1, Property 2 is obtained.
@@ 5.1 Model
Applying Riemann transformations ${z}(x,t)={u_{t}}(x,t)-\sqrt{\frac{T_{0}}{\rho}}{u_{x}}(x,t)$, ${w}(x,t)={u_{t}}(x,t)+\sqrt{\frac{T_{0}}{\rho}}{u_{x}}(x,t)$ and defining a new variable $Y(t)=[y_{1}(t),\dot{y}_{1}(t)]^{T}$, (143)–(145) can be rewritten as
@@
where $q_{1}=q_{2}=\sqrt{\frac{T_{0}}{\rho}},c=\frac{{{-d_{c}}}}{2\rho},p=q=1$, $A=\left[0,1;0,\frac{-d_{0}}{M_{0}}-\frac{\sqrt{T_{0}\rho}}{M_{0}}\right],B=[0,\frac{\sqrt{T_{0}\rho}}{M_{0}}]^{T},C=[0,2].$ After the Riemann transformation in Sec. 5.1, the airflow disturbances $f_{i}$ in the wave PDEs is given by the external disturbances $d(t)=\bar{P}_{d}v_{d}(t)$ (14) with $\bar{P}_{d}=I_{4\times 4}$ and $v_{d}=[\sin(0.25t),\cos(0.25t),\sin(0.5t),\cos(0.5t)]$ in the $2\times 2$ hyperbolic PDEs (146)–(150), where $G_{1}=[g_{1},g_{2}]^{T}$, $g_{1}=[1,0,0,0]$, $g_{2}=[1,1,1,1]$, $G_{4}=[0,1,0,1]$, $G_{5}=[1,0,1,0]$, $G_{2}(x)=[x,0,0,0]$, $G_{2}(x)=[0,x,0,0]$. The desired reference is set as $r(t)=\sin(0.25\pi t)+\cos(0.25\pi t)$, i.e., $\bar{P}_{r}=[1,1]$, and $v_{r}=[\sin(0.25\pi t),\cos(0.25\pi t)]$ in (15). Therefore, the system matrix of exogenous signal model is $S_{d}=[0,0.25,0,0;-0.25,0,0,0;0,0,0,0.5;0,0,$ 0.5,0]-0.5,0] and $S_{r}=[0,0.25\pi;-0.25\pi,0]$.
@@
Compared to the plant described by (1)–(5), the simulation model represented by (146)–(150) includes additional terms, specifically $c_{1}z(x,t)$ and $c_{2}w(x,t)$ in (148) and (149), respectively. However, these additions do not alter the state-feedback control design presented in this paper. To be exact, applying state-feedback control, the target system defined by (68) and (70), which corresponds to the closed-loop system, remains unchanged except for the updates that (69) are modified to $\beta_{t}=q_{2}\beta_{x}+c\beta$ and $\alpha_{t}=-q_{1}\alpha_{x}+c\alpha$. These modifications clearly do not impact the safety and stability results obtained. The only change in the observer-based output-feedback controller is adding the terms $c\hat{z}(x,t)$ and $c_{2}\hat{w}(x,t)$ in (90) and (91) of the observer, while all other aspects remain unchanged.
The initial values are defined as $w(x,0)=\cos(2\pi x)$, $z(x,0)=\sin(3\pi x)$, $y_{2}(0)=0$, $\hat{w}(x,0)=\cos(2\pi x)+0.5$, $\hat{z}(x,0)=\sin(3\pi x)+0.5$, $\hat{Y}(0)=[\hat{y}_{1}(0),\hat{y}_{2}(0)]^{T}=[0,0]^{T}$, $\hat{v}_{d}(0)=[0.5,1.5,0.5,1.5]^{T}$, $\hat{v}_{r}(0)=[0.5,1.5]^{T}$ where the initial estimation errors are $0.5$. Especially, the values $y_{1}(t_{0})$ will be given in the next subsections. The simulation is conducted using the finite difference method with the space step as $0.05$ and the time space $0.001$.
@@ 5.2 Controller and simulation results
We test our safe controller in the following two cases.
5.2.1 Case 1. The safe region is on one side of the desired reference.
The barrier function is set as $h(e(t),t)=e(t)=y_{1}(t)-r(t)$, where $r(t)$ is the desired reference of UAV, and where the region above $r(t)$ is safe. Then we know $\vartheta=\frac{\partial h(e(t),t)}{\partial e}=1$. According to (47)–(49), we have $h_{1}=z_{1}(t)+\sigma(t)$, $h_{2}=z_{2}(t)+\dot{\sigma}(t)+k_{1}z_{1}(t)+k_{1}\sigma(t)$, where $\sigma(t)$ is given by (49) with choosing $\epsilon=2$ and $t_{a}=2$. According to (54), $f(\underline{z}_{n}(t),t)$ is obtained as, $f(\underline{z}_{n}(t),t)=\frac{1}{b}((k_{1}+k_{2})z_{2}+k_{1}k_{2}z_{1}+\ddot{\sigma}(t)+(k_{1}+k_{2})\dot{\sigma}(t)+k_{1}k_{2}\sigma(t))$. Then $\varsigma(1,t)$ can be obtained by recalling (72) that uses (B.5). Then the state-feedback controller is obtained recalling (71), where the design parameters are chosen as $k_{1}=1.5$, $k_{2}=4$, and the output-feedback controller $U_{f}$ is obtained applying (137) with choosing the control design parameters $k_{1}=3$, $k_{2}=4$, $M_{c}=15$, $\sigma_{r}=2$, and the observer design parameters $L_{d}=[11,8,11,6]$, $L_{r}=[2,1]$. For safe initialization, we set $y_{1}(t_{0})=8$. For the unsafe initialization case, we set $y_{1}(t_{0})=-1$ that is beyond the safe region, and the design parameters in the output-feedback control law are the same as the output-feedback control in the safe initialization except for $k_{1},k_{2}$ that are chosen as $k_{1}=2.5$, $k_{2}=4.5$ now, and additionally choosing $t_{a}=2$, $\epsilon=2$ in (49) to enforce the out return to the safe region. The simulation results are presented below.
@@
(a) Estimates of $\hat{v}_{d}$ and $\hat{v}_{r}$
(b) Control inputs
Figure 6: Estimates of exogenous signals $v_{d}$, $v_{r}$ given in Sec. 5.1 (red and blue lines are estimates $\hat{v}_{d}$ and $\hat{v}_{r}$, respectively, and dashed black lines are their true values) as well as the control inputs.
We know from the open-loop responses of $y_{1}$ in Fig. 2 that the plant is open-loop unstable, while both state-feedback control and output-feedback control achieve effective safe regulation— tracking the reference and ensuring the position of the payload stays within the safe region. Moreover, when $y_{1}$ starts outside the safe region, the designed control input effectively guides it back to safety within 1.85 seconds, which is less than the designated time $\bar{t}_{0}+t_{a}=2.058$ in the control design. Besides, we know from Fig. 2 that observer state $\hat{y}_{1}$ fast converges to the true value, i.e., the blue dashed line that denotes the evolution of $y_{1}$, whose initial state is in the safe region, under the output-feedback. Another ODE state $y_{2}$, i.e., the payload velocity, is shown in Fig. 3, and the PDE states that reflect the oscillations along the cable are given in Fig. 4 (where we only show the case that the output-feedback control from safe initial states for avoiding repetition), demonstrating the boundedness of these signals. It is also shown in Fig. 3 that the observer state $\hat{y}_{2}$ rapidly converges to its true value $y_{2}$, denoted by the blue dashed line, in the same case (safe initialization). In addition to ODE state estimates, the observer errors for PDE states and the exogenous signals can be seen in Figs 5, 6(a), which show that the observer can successfully track the unmeasured PDE states and the exogenous signals $v_{d}$ and $v_{r}$. The control inputs in the above three situations are given in Fig. 6(b).
@@
Because $h$ is not differentiable at $e=0$, the condition in Assumption 7 is not strictly satisfied by (151). However, the controller remains valid by explicitly defining $\frac{\partial h(e,t)}{\partial e}=-1$ when $e=0$, i.e., $\vartheta(e(t),t)=-1$ for $e(t)>0$ and $\vartheta(e(t),t)=1$ for $e(t)\leq 0$. The signal $\vartheta$ is now discontinuous at $e=0$, violating the continuity requirement in Assumption 7. This implies that ${\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))$ in the output-feedback controller is not guaranteed to be invariant as in the proof of Theorem 2. Nevertheless, with appropriate observer gains $L_{d}$, $L_{r}$ chosen via pole placement, the observer errors can rapidly converge to zero, which means that $\rho_{e}(t)$ decays quickly to zero and the influence of the term ${\rm sign}(\vartheta(z_{1}(t_{0}),t_{0}))\rho_{e}$ in the output-feedback controller becomes negligible after a short initial period. Therefore, the proposed control design remains valid without modification, as shown in the upcoming simulation results.
According to (47)–(49), we have $h_{1}=-|z_{1}(t)|+\delta(t)+\sigma(t)$, $h_{2}=\vartheta z_{2}+\dot{\delta}(t)+\dot{\sigma}(t)-k_{1}|z_{1}(t)|+k_{1}\delta(t)+k_{1}\sigma(t)$, with $\sigma(t)$ is given by (49) with choosing $\epsilon=2$ and $t_{a}=2$. Recalling (54), we obtain $f$ as $f(\underline{z}_{n}(t),t)=\frac{1}{b}\big((k_{1}+k_{2})\vartheta z_{2}-k_{1}k_{2}|z_{1}(t)|+\ddot{\delta}(t)+(k_{1}+k_{2})\dot{\delta}(t)+k_{1}k_{2}\delta(t)+\ddot{\sigma}(t)+(k_{1}+k_{2})\dot{\sigma}(t)+k_{1}k_{2}\sigma(t)\big)$. Then $\varsigma(1,t)$ can be obtained by recalling (72) that uses (B.5). The output-feedback controller $U_{f}$ is obtained applying (137), (138), where all control design parameters are the same with those in cases 1 except for $k_{1}=30$, $k_{2}=16$, $M_{c}=180$, $\sigma_{r}=1$. For safe initialization, we set $y_{1}(t_{0})=-10$. For unsafe initialization, we set $y_{1}(t_{0})=20$, which falls outside the safe region. The values of control design parameters remain the same in both situations of safe initialization and unsafe initialization. Additionally, we choose $t_{a}=3$ and $\epsilon=2$ in (49) to ensure that the system’s output returns to the safe region in the unsafe initialization situation. The simulation results are presented below.
@@
It is observed from Fig 7 that when $y_{1}$, i.e., the position of the payload, starts within the safe region, it successfully tracks the reference meanwhile avoiding collisions with the barriers all the time; when the $y_{1}$ starts outside the safe region, it will return to the safe region at 0.34 s, which is much less than $\bar{t}_{0}+t_{a}=2.058$, and also eventually follow the reference. Fig. 8 shows the PDE states are bounded even though anti-damping is added to challenge the task, so is $y_{2}$ whose result is not presented here due to space limit. The results about the observer errors and the estimation errors of the external signals are similar to Figs. 5 and 6, which are not present here for avoiding repetition.